OpenStudy (anonymous):

L1:x=−2t,y=1+2t,z=3t and L2:x=−9+5s,y=4+1s,z=3+3s Find the point of intersection of the two lines.

ganeshie8 (ganeshie8):

set the components equal to each other and solve s, t

OpenStudy (anonymous):

-2t = -9 + 5s ? like that?

ganeshie8 (ganeshie8):

Yes! -2t = -9+5s 1+2t = 4+1s 3t = 3+3s the lines will intersect if there is a solution to this system of 3 equations with 2 unknowns

OpenStudy (anonymous):

OOH! that makes sense. Similar to finding intersection in 2-space.

OpenStudy (anonymous):

going to try it out, thank you @ganeshie8

ganeshie8 (ganeshie8):

np :) you may simply pick any two equations and solve s,t once you have s,t values plug them in the third equation and check if they satisfy

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

woo got it, ty :)